OpenStudy (anonymous):

4x+6y=8 GIVE ANY NUMBER FOR x and y. ( from -10 to 10) other than x=2 and y=0

OpenStudy (agreene):

If you solve this to a y=x equation, you can plug in values of x and find the respective values of y. so, throw your 4x on the other side: 6y=8-4x then divide by 6 and you have y=(8-4x)/6 which you can factor to y= (-2(x-2)/3 now you can plug in values of x from -10 to 10 and get the y values that correspond.